Please HELP> i am not good with fractions.
Solve by completing the square
m^2-1/2m=7/2

Answer by stanbon(75887)

Solve by completing the square
m^2-1/2m=7/2
----------------
m^2 -(1/2)m + (1/4)^2 = (7/2) + (1/4)^2
(m - (1/4))^2 = 56/16+1/16
(m - (1/4)) = +-(1/4)sqrt(57)
m = (1/4)+(1/4)sqrt(57) or m = (1/4)-(1/4)sqrt(57)
=========================================================================
Cheers,
Stan H.